OpenStudy (amber15):

The intensity I of light bulb measured in watts per square meter (w/m^2), varies inversely as the square of the distance d from the light bulb . . suppose I is 45 w/m^2 when the distance is 5 m . find the intensity when the distance from the light bulb is 7 m away

jimthompson5910 (jim_thompson5910):

"The intensity I of light bulb...varies inversely as the square of the distance d from the light bulb" so I = k/(d^2)

jimthompson5910 (jim_thompson5910):

Plug in I = 45 and d = 5 I = k/(d^2) 45 = k/(5^2) k = ???

OpenStudy (amber15):

would it be 1125/49

OpenStudy (amber15):

i figured it over hours of trying

jimthompson5910 (jim_thompson5910):

that is correct

jimthompson5910 (jim_thompson5910):

I = 1125/49 when d = 7

OpenStudy (amber15):

awesome but i will be back cause i learn best from you